# First Measurement of Solar 8B Neutrinos via Coherent Elastic Neutrino-Nucleus Scattering with XENONnT

### v1

@@ Abstract
Fig. 4 shows the first measurement of the flux-weighted CENS cross-section on Xe as , consistent with the SM prediction of 1.2. Since the momentum transferred from a solar neutrino to a Xe nucleus is /c, this measurement is less sensitive to uncertainties in the nuclear form factor compared to CENS measurements made by the COHERENT collaboration with neutrinos produced by the SNS [42]. The measurements of the flux-weighted CENS cross-section on CsI [42], Ar [4], and Ge [43] nuclei by the COHERENT Collaboration are shown in Fig. 4 for comparison. Due to the lower average energy, solar neutrino flux-weighted CENS cross-section is the lowest one measured to date.
Figure 4: Measurements of the flux-weighted CENS cross-section . The measurement using Xe nuclei solar neutrinos from this work is shown in black. The 90% CL upper limit from XENON1T [6] is shown in blue. The measurements with neutrinos from the SNS by the COHERENT collaboration using CsI [42] (red), Ar [4] (green) and Ge [43] (orange) nuclei are also shown. For comparison, the SM predictions are shown by vertical dashed lines, assuming the nuclear form factors from [44].
Summary — We performed a blind search for NR signals from solar neutrinos via CENS with XENONnT using data from two science runs with a combined exposure of 3.51 ty. By lowering the S1 and S2 thresholds, we are able to include NR signals as low as 0.5 keV. Various techniques are developed to reduce the dominant AC background. Various calibrations, including and 37Ar, are performed to understand the detector response, signal, and background modeling. The data disfavors the background-only hypothesis at 2.73 . The unconstrained best-fit number of CENS signals is , consistent with the expectation of events, based on the measured solar neutrino flux from SNO [22], the theoretical CENS cross-section with Xe nuclei [23], and the calibrated detector response to low-energy NRs in XENONnT. Thus the measured solar neutrino flux is , consistent with SNO, and measured neutrino flux-weighted CENS cross-section on Xe is , consistent with the SM prediction. As XENONnT continues to take data, more precise measurements are expected in the future.
@@
We would like to thank the COHERENT Collaboration for providing data points and predictions for the measurement of flux-weighted CENS cross-section at the SNS.
@@
The S2 thresholds for both SR0 and SR1 are thus conservatively increased for the CENS search data, with minor loss in the discovery potential of the solar CENS signals. The final prediction of the AC background and observations in the AC sideband are shown in Tab. 2. In SR1, 326 events are observed with an expectation of 302.5, yielding a p-value of 0.16. The projection of the 4 analysis dimensions with the same binning used in CENS search in sideband data with the larger than in both SR0 and SR1 data are shown in Fig. 5.
@@
The signal and background modeling is validated by the measurement of of 37Ar L-shell EC, which is performed with a blind analysis. The AC background in this measurement is estimated to be 110955 based on a similar modeling approach to that in the CENS search. After unblinding, 1676 events are observed. The observed events above the expected AC background are strongly validated by a four-dimensional GOF test, yielding a p-value of 0.72. Fig. 6 shows the observed events in the same analysis dimensions as the CENS search along with the AC background and the best-fit 37Ar L-shell EC signal during the 37Ar calibration. The measurement will be presented in a future publication.

### v2

@@ Abstract
Fig. 4 shows the first measurement of the flux-weighted CENS cross section on Xe as , consistent with the SM prediction of 1.2. Since the momentum transferred from a solar neutrino to a Xe nucleus is /c, this measurement is less sensitive to uncertainties in the nuclear form factor compared to CENS measurements made by the COHERENT collaboration with neutrinos produced by the SNS [44]. The measurements of the flux-weighted CENS cross section on CsI [44], Ar [5], and Ge [45] nuclei by the COHERENT Collaboration are shown in Fig. 4 for comparison. Because of the lower average energy, the solar neutrino flux-weighted CENS cross section is the lowest one measured to date.
Figure 4: Measurements of the flux-weighted CENS cross section . The measurement using Xe nuclei solar neutrinos from this work is shown in black. The 90% CL upper limit from XENON1T [7] is shown in blue. The measurements with neutrinos from the SNS by the COHERENT collaboration using CsI [44] (red), Ar [5] (green) and Ge [45] (orange) nuclei are also shown. For comparison, the SM predictions are shown by vertical dashed lines.
Summary — We performed a blind search for NR signals from solar neutrinos via CENS with XENONnT using data from two science runs with a combined exposure of 3.51 tyr. By lowering the S1 and S2 thresholds, we are able to include NR signals as low as 0.5 keV. Various techniques are developed to reduce the dominant AC background. Various calibrations, including and 37Ar, are performed to understand the detector response, signal, and background modeling. The data disfavor the background-only hypothesis at 2.73 . The unconstrained best-fit number of CENS signals is , consistent with the expectation of events, based on the measured solar neutrino flux from SNO [23], the theoretical CENS cross section with Xe nuclei [25], and the calibrated detector response to low-energy NRs in XENONnT. Thus, the measured solar neutrino flux is , consistent with SNO, and the measured neutrino flux-weighted CENS cross section on Xe is , consistent with the SM prediction. As XENONnT continues to take data, more precise measurements are expected in the future.
@@
We would like to thank the COHERENT Collaboration for providing data points and predictions for the measurement of flux-weighted CENS cross section at the SNS.
@@
Note Added —
Recently, we noticed the results of the neutrino flux measurement from the PandaX Collaboration with a similar statistical significance in [46].
@@
The S2 thresholds for both SR0 and SR1 are thus conservatively increased for the CENS search data, with minor loss in the discovery potential of the solar CENS signals. The final prediction of the AC background and observations in the AC sideband are shown in Tab. 2. The projection of the four analysis dimensions with the same binning used in the CENS search in sideband data with the larger than in both SR0 and SR1 data are shown in Fig. 5.
@@
The signal and background modeling is validated by the measurement of of 37Ar L-shell EC, which is performed with a blind analysis. The AC background in this measurement is estimated to be 106253 based on a similar modeling approach to that in the CENS search. After unblinding, 1676 events are observed. The observed events above the expected AC background are strongly validated by a four-dimensional GOF test, yielding a p-value of 0.92. Fig. 6 shows the observed events in the same analysis dimensions as the CENS search along with the AC background and the best-fit 37Ar L-shell EC signal during the 37Ar calibration. The measurement will be presented in a future publication.
Appendix C: Separate SR0/SR1 best-fit results —
The distributions of the observed events in SR0 and SR1 and the corresponding best-fit model projected to each analysis dimension are shown individually in Fig. 7.
Figure 7: Distributions of best-fit signal and background, together with the data in the projected analysis dimensions, with SR0 and SR1 shown in the left and right column, respectively. The observed number of events with Poisson uncertainties in each bin is shown in black. The CENS signal is represented by the light green histogram on top of the backgrounds, which are indicated by purple (AC), blue (ER), and yellow (neutron) histograms.
Kopeliovich and Frankfurt [1974]V. B. Kopeliovich and L. L. Frankfurt, Isotopic and chiral
structure of neutral current, JETP Lett. 19, 145 (1974).
Drukier and Stodolsky [1984]A. Drukier and L. Stodolsky, Principles and
Applications of a Neutral Current Detector for Neutrino Physics and
Astronomy, Phys. Rev. D 30, 2295 (1984).
Billard et al. [2014]J. Billard, L. Strigari, and E. Figueroa-Feliciano, Implication
of neutrino backgrounds on the reach of next generation dark matter direct
detection experiments, Phys. Rev. D 89, 023524 (2014), arXiv:1307.5458 [hep-ph]
Aprile et al. [2024c]E. Aprile et al. (XENON), XENONnT Analysis: Signal Reconstruction,
Calibration and Event Selection, (2024c), arXiv:2409.08778 [hep-ex]
Bahcall et al. [1996]J. N. Bahcall, E. Lisi,
D. E. Alburger, L. De Braeckeleer, S. J. Freedman, and J. Napolitano, Standard neutrino spectrum from B-8 decay, Phys. Rev. C 54, 411 (1996), arXiv:nucl-th/9601044 .
Barranco et al. [2005]J. Barranco, O. G. Miranda, and T. I. Rashba, Probing new physics with
coherent neutrino scattering off nuclei, JHEP 12, 021, arXiv:hep-ph/0508299 .
Aprile et al. [2024d]E. Aprile et al. (XENON), XENONnT WIMP Search: Signal & Background
Modeling and Statistical Inference, (2024d), arXiv:2406.13638
[physics.data-an] .
Aprile et al. [tion]E. Aprile et al. (XENON), (In preparation), .
Aprile et al. [2006]E. Aprile, C. E. Dahl,
L. DeViveiros, R. Gaitskell, K. L. Giboni, J. Kwong, P. Majewski, K. Ni, T. Shutt, and M. Yamashita, Simultaneous
measurement of ionization and scintillation from nuclear recoils in liquid
xenon as target for a dark matter experiment, Phys. Rev. Lett. 97, 081302 (2006), arXiv:astro-ph/0601552
Kopec et al. [2021]A. Kopec, A. L. Baxter,
M. Clark, R. F. Lang, S. Li, J. Qin, and R. Singh, Correlated single- and few-electron backgrounds milliseconds after
interactions in dual-phase liquid xenon time projection chambers, JINST 16 (07), P07014, arXiv:2103.05077 [physics.ins-det]
Sorensen [2017]P. Sorensen, Electron train
backgrounds in liquid xenon dark matter search detectors are indeed due to
thermalization and trapping, (2017), arXiv:1702.04805 [physics.ins-det]
Szydagis et al. [2022]M. Szydagis et al., A Review
of NEST Models, and Their Application to Improvement of Particle
Identification in Liquid Xenon Experiments, (2022), arXiv:2211.10726 [hep-ex] .
Adamski et al. [2024]S. Adamski et al., First
detection of coherent elastic neutrino-nucleus scattering on germanium, (2024), arXiv:2406.13806 [hep-ex] .